# Quantum circuit generation via test-time learning with large language models

### v1

@@ Abstract
Large language models (LLMs) can generate structured artifacts, but turning them into dependable optimizers for scientific design requires closed-loop refinement under black-box evaluation. We formulate quantum circuits synthesis as a test-time optimization problem in which an LLM proposes edits to a fixed-length gate list and an external simulator scores the resulting state using the Meyer-Wallach (MW) global entanglement measure. We introduce a lightweight test-time learning recipe that (i) reuses previously high-scoring candidates as an explicit memory trace, (ii) augment prompts with score-difference feedback, and (iii) employs restart-from-best sampling to escape plateaus. In fixed 20-qubit settings, demonstrates the feasibility of the approach but with a low success rate. Then, we move on 25 qubit, adding feedback. The full strategy mitigates a pronounced plateau observed under naive querying. Beyond raw scores, we analyze the structure of the synthesized states and find that high-MW solutions often resemble stabilizer or graph-state-like constructions, while full connectivity is not guaranteed-reflecting both properties of the MW metric and the constraints induced by prompt design. Overall, our results demonstrate the promise-and the limitations- of memory- and evaluator-guided LLM optimization for quantum circuit synthesis, and underscore the continued importance of careful objective selection and human-in-the-loop problem formulation.
@@ I Introduction
In this work we explore the potential of test-learning approach, avoiding completely the need to prepare a dataset for training or fine-tuning the system. By casting the circuit synthesis as a test-time optimization loop in which an LLM proposes circuit edits from a single initially random circuit. In this way, the generation process becomes a trajectory that can accumulate “skills” specific to the task, such as reusable entangling motifs and topology patterns. We represent circuits directly as gate lists, similarly to other recent job that focus on pre-trained or fine-tuning approaches [llm6, 29], and restrict available operations to an experimentally motivated discrete subset, avoiding continuous parameters that would inflate the search space. As the validation metric, we opt for the Meyer–Wallach global entanglement measure, an experimentally-friendly and permutation-invariant metric defined in terms of single-qubit reduced purities. This choice deliberately emphasizes global structure rather than local heuristics: the model is rewarded only when the entire register becomes strongly entangled.
@@ III Main Results
Here, we explore the application of GPT5.1 with a memory step and no feedback. We experiment using 25 to 45 gates on 20-qubit circuits. This choice shrinks the search space, but dodges a trick recurrently applied by small models, which is the application of a high number ( num gates) of random CNOTs. Ranging from 25 to 45 gates, we can see from Table 1 that the model improves over the initial random circuit, and can obtain up to perfect MW in some situations, as we can see for 35 gates. Nevertheless, the success probability of having an MW measure is . This evidence backs up our insights, demonstrating that LLMs can sample optimal circuits beyond repeated sampling approaches, also known as the Large Language Monkeys [brown2024largelanguagemonkeysscaling]. Nevertheless, to achieve greater performances, a leap is required.
@@ IV Conclusions
In this work, we have analyzed the ability of a new approach to quantum circuit synthesis using large language models for a global entanglement metric. We demonstrate that thinking models combined with test-time inference and feedback can obtain high values of global entanglement without any post-training. Firstly, we have designed a query with and without feedback and test its efficacy for a single run, obtaining outcomes of the order of 0.99 for 20 qubits and 35 gates, when the Meyer-Wallach metric is used to evaluate them. When moving to 25 qubits, we implement a GPT5.2 model plugging a feedback inside the prompt, and running each experiment 3 times in a row, re-using the best previous outcome to introduce a degree of perturbation. This approach immediately steers the outcomes toward peaks of MW in the synthesized solution, in a single run. This demonstrates the pivotal role played by feedback and simple memory usage for circuit synthesis with LLM, paving the way to new approaches in circuit synthesis. Interestingly, we also found that some circuits behave like ”local minima”, stopping the model from sampling improvement. This is likely due to an intrinsic limitation in the model’s reasoning ability, and is likely spawned by the lack of high-quality pre-selected theoretical knowledge, which can be used to fine-tune the model to achieve a higher level of performance and solution quality.

### v2

@@ Abstract
Large language models (LLMs) can generate structured artifacts, but turning them into dependable optimizers for scientific design requires closed-loop refinement under black-box evaluation. We formulate quantum circuits synthesis as a test-time optimization problem in which an LLM proposes edits to a fixed-length gate list and an external simulator scores the resulting state using the Meyer-Wallach (MW) global entanglement measure. We introduce a lightweight test-time learning recipe that (i) reuses previously high-scoring candidates as an explicit memory trace, (ii) augments prompts with score-difference feedback, and (iii) employs restart-from-best sampling to escape plateaus. In fixed 20-qubit settings, it demonstrates the feasibility of the approach, but with a low success rate. Then, we move on to 25 qubits, adding feedback. The full strategy mitigates a pronounced plateau observed under naive querying. Beyond raw scores, we analyze the structure of the synthesized states and find that high-MW solutions often resemble stabilizer or graph-state-like constructions, while full connectivity is not guaranteed, reflecting both properties of the MW metric and the constraints induced by prompt design. Overall, our results demonstrate the promise-and the limitations- of memory- and evaluator-guided LLM optimization for quantum circuit synthesis, and underscore the continued importance of careful objective selection and human-in-the-loop problem formulation.
@@ I Introduction
In this work, we explore the potential of a test-learning approach, avoiding the need to prepare a dataset for training or fine-tuning the system. By casting the circuit synthesis as a test-time optimization loop in which an LLM proposes circuit edits from a single initially random circuit. In this way, the generation process becomes a trajectory that can accumulate “skills” specific to the task, such as reusable entangling motifs and topology patterns. We represent circuits directly as gate lists, similarly to other recent jobs that focus on pre-trained or fine-tuning approaches [llm6, 29], and restrict available operations to an experimentally motivated discrete subset, avoiding continuous parameters that would inflate the search space. As the validation metric, we opt for the Meyer–Wallach global entanglement measure, an experimentally-friendly and permutation-invariant metric defined in terms of single-qubit reduced purities. This choice deliberately emphasizes global structure rather than local heuristics: the model is rewarded only when the entire register becomes strongly entangled.
@@ III Main Results
Here, we explore the application of GPT5.1 with an episodic memory step and no feedback. We experiment using 25 to 45 gates on 20-qubit circuits. This choice shrinks the search space, but dodges a trick recurrently applied by small models, which is the application of a high number ( num gates) of random CNOTs. Ranging from 25 to 45 gates, we can see from Table 1 that the model improves over the initial random circuit, and can obtain up to perfect MW in some situations, as we can see for 35 gates. Nevertheless, the success probability of having an MW measure is . This evidence backs up our insights, demonstrating that LLMs can sample optimal circuits beyond repeated sampling approaches, also known as the Large Language Monkeys [brown2024largelanguagemonkeysscaling]. Nevertheless, to achieve greater performances, a leap is required.
@@ IV Conclusions
In this work, we have analyzed the ability of a new approach to quantum circuit synthesis using large language models for a global entanglement metric. We demonstrate that thinking models combined with test-time inference and feedback can obtain high values of global entanglement without any post-training. Firstly, we have designed a query with and without feedback and test its efficacy for a single run, obtaining outcomes of the order of 0.99 for 20 qubits and 35 gates, when the Meyer-Wallach metric is used to evaluate them. When moving to 25 qubits, we implement a GPT5.2 model plugging a feedback inside the prompt, and running each experiment 3 times in a row, re-using the best previous outcome to introduce a degree of perturbation. This approach immediately steers the outcomes toward peaks of MW in the synthesized solution, in a single run. This demonstrates the pivotal role played by feedback and a simple episodic memory usage for circuit synthesis with LLM, paving the way to new approaches in circuit synthesis. Interestingly, we also found that some circuits behave like ”local minima”, stopping the model from sampling improvement. This is likely due to an intrinsic limitation in the model’s reasoning ability, and is likely spawned by the lack of high-quality pre-selected theoretical knowledge, which can be used to fine-tune the model to achieve a higher level of performance and solution quality.